Royal Gift Shop sells cards, supplies, and various holiday greeting cards. Sales to retail customers are subject to an 8 percRequired: 2. Record the transactions in a general journal. 3. Post the entries from the general journal to the appropriate acGeneral Journal General Ledgers Analyze Record the transactions in a general journal. (Round your intermediate calculations aGeneral Journal General Ledgers Analyze Post the entries from the general journal to the appropriate accounts in the generalGeneral Journal General Ledgers Analyze What was the total credit card expense incurred in February? (Round your intermediate

Need help please asap.
Answer #1

1. Recording of Journals :

Date Account Title Debit Credit
01 February 2019 Accounts Receivable Dr. 3290
To Sales 3290
(Being sold crystal goods to Lovely Kitchens.- Invoice:5950)
15 February 2019 Cash Dr. 9936
To Sales 9200
To Sales Tax payable 736
(Being Cash Sales from 01/02 to 15/02)
15 February 2019 Receivable from Credit Card Dr 13716
To Sales 12700
To Sales Tax payable 1016
(Being Credit Card sales from 01/02 to 15/02)
Credit Card Expense Dr 137.16
To Receivable from Credit Card 137.16
(Being Credit card expense from 01/02 to 15/02 @1%)
Cash Dr 13578.84
To Receivable from Credit Card 13578.84
(Being received net amount receivable from Credit Card Companies.)
16 February 2019 Cash Dr 3290
To accounts Receivable 3290
(Being Cheque received from Lovely Kitchens)
16 February 2019 Receivable fromAmerican Express Dr 9396
To Sales 8700
To Sales Tax payable 696
(Being merchandise sold through American Express card.)
17 February 2019 Accounts Receivable Dr. 6960
To Sales 6960
(Being sold Roman Statues to Beautiful Bedrooms.- Invoice:5951)
20 February 2019 Cash Dr 9208.08
American Express Credit card expense 187.92
To Receivable fromAmerican Express 9396
(Being American Express Credit card expense @2%)
27 February 2019 Cash Dr 6960
To accounts Receivable 6960
(Being Cheque received from Beautiful Bedrooms)
28 February 2019 Cash Dr. 8208
To Sales 7600
To Sales Tax payable 608
(Being Cash Sales from 16/02 to 28/02)
28 February 2019 Receivable from Credit Card Dr 16200
To Sales 15000
To Sales Tax payable 1200
(Being Credit Card sales from 16/02 to 28/02)
Credit Card Expense Dr 162
To Receivable from Credit Card 162
(Being Credit card expense from 16/02 to 28/02 @1%)
Cash Dr 16038
To Receivable from Credit Card 16038
(Being received net amount receivable from Credit Card Companies.)
28 February 2019 Receivable fromAmerican Express Dr 10692
To Sales 9900
To Sales Tax payable 792
(Being merchandise sold through American Express card.)

2. Posting of entries to Ledgers :

Cash Account No. 101 Accounts Receivable Account No. 121
Date Debit Credit Balance Date Debit Credit Balance
01-Feb 22930 0 22930 01-Feb 3290 0 3290
15-Feb 9936 0 32866 16-Feb 0 3290 0
15-Feb 13578.84 0 46444.84 17-Feb 6960 0 6960
16-Feb 3290 0 49734.84 27-Feb 0 6960 0
20-Feb 9208.08 0 58942.92
27-Feb 6960 0 65902.92
28-Feb 8208 0 74110.92
28-Feb 16038 0 90148.92
Sales Tax Payable Account No. 222 Sales Account No. 401
Date Debit Credit Balance Date Debit Credit Balance
15-Feb 0 736 736 01-Feb 0 3290 3290
15-Feb 0 1016 1752 15-Feb 0 9200 12490
16-Feb 0 696 2448 15-Feb 0 12700 25190
28-Feb 0 1200 3648 16-Feb 0 8700 33890
28-Feb 0 792 4440 17-Feb 0 6960 40850
28-Feb 0 15000 55850
28-Feb 0 9900 65750
Credit card Expense Account No. 521
Date Debit Credit Balance
15-Feb 137.16 0 137.16
20-Feb 187.92 0 325.08
28-Feb 162 0 487.08

3. Total Credit card expense incurred in February :

Credit card expense as per ledger = 487.08

+ Expense for sales through American Express card on 28/02 = 10692*2%=213.84

Total Credit card Expense = 487.08 + 213.84 = $700.92

(If found helpful,kindly rate positively.)